A company is considering two mutually exclusive expansion plans. Plan A requires a $39 million expenditure on a large-scale integrated plant that would provide expected cash flows of $6.23 million per year for 20 years. Plan B requires a $11 million expenditure to build a somewhat less efficient, more labor-intensive plant with an expected cash flow of $2.47 million per year for 20 years. The firm's WACC is 11%. The data has been collected in the Microsoft Excel Online file below. Open the spreadsheet and perform the required analysis to answer the questions below.

- Calculate each project's NPV. Round your answers to two decimal places. Do not round your intermediate calculations. Enter your answers in millions. For example, an answer of $10,550,000 should be entered as 10.55.

Plan A: $   million

Plan B: $   million

Calculate each project's IRR. Round your answer to two decimal places.

Plan A: __%

Plan B: __%

- By graphing the NPV profiles for Plan A and Plan B, approximate the crossover rate to the nearest percent.

__%

- Calculate the crossover rate where the two projects' NPVs are equal. Round your answer to two decimal places.

__%

- Why is NPV better than IRR for making capital budgeting decisions that add to shareholder value?
Answer #1

1.
=-39+6.23/11%*(1-1/1.11^20)=10.6115341711957

2.
=-11+2.47/11%*(1-1/1.11^20)=8.6694204498962

3.
=RATE(20,-6.23,39)=14.9979778746622%

4.
=RATE(20,-2.47,11)=22.0362%

5.
13%

6.
=RATE(20,-6.43+2.47,39-11)=12.8917553881267%

7.
Consistent with shareholder wealth maximization, gives value in absolute terms of value added, gives only one value